Reduce the quadratic form X2+3Y2+3Z2-2yz to the canonical form through orthogonal transformation

1
Expert's answer
2022-01-18T07:16:56-0500

We shall reduce the quadratic form "x^{2}+3y^{2}+3z^{2}-2yz" to the canonical form through the orthogonal transformation.


NOTE: WE SHALL RE-REPRESENT THE VARAIBALE x, y and z TO VARIABLES "x_{1},x_{2}\hspace{0.1cm}and\hspace{0.1cm}x_{3}" respectively.


NOW, using "x_{1},x_{2}\hspace{0.1cm}and\hspace{0.1cm}x_{3}"

"\displaystyle Q(x)=x_{1}^{2}+3x_{2}^{2}+3x_{3}^{2}-2x_{2}x_{3}"

The symmetric matrix A with the given quadratic form is

"A=" "\begin{bmatrix}
 1 & 0 & 0 \\
 0 & 3 & -1\\
 0 & -1 & 3 
\end{bmatrix}"

The characteristic equation of A is

"\mid(A-\lambda I)\mid=\lambda^{3}-7\lambda^{2}+14\lambda-8=0\\\implies \lambda=1,2,4"


The corresponding eigen values are:

for "\lambda_{1}=1,\hspace{0.2cm}is\hspace{0.2cm}X_{1}=k\begin{bmatrix}
 1 \\
 0\\
 0 
\end{bmatrix}"


for "\lambda_{2}=2,\hspace{0.2cm}is\hspace{0.2cm}X_{2}=k\begin{bmatrix}
 0 \\
 1\\
 1
\end{bmatrix}"


for "\lambda_{3}=1,\hspace{0.2cm}is\hspace{0.2cm}X_{3}=k\begin{bmatrix}
 0 \\
 1\\
 -1
\end{bmatrix}"


Clearly, the eigen vectors "X_{1}=\begin{bmatrix}
 1 \\
 0\\
 0 
\end{bmatrix}" , "X_{2}=\begin{bmatrix}
 0 \\
 1\\
 1
\end{bmatrix}" and "X_{3}=\begin{bmatrix}
 0 \\
 1\\
 -1
\end{bmatrix}" are linearly independent and orthogonal.


The normalized vectors of the above eigen vextors "X_{1}, X_{2},X_{3}" are:

"e_{1}=\frac{X_{1}}{\mid\mid X_{1}\mid\mid}= \begin{bmatrix}
 1 \\
 0\\
 0 
\end{bmatrix}" , "\displaystyle e_{2}=\frac{X_{2}}{\mid\mid X_{2}\mid\mid}= \begin{bmatrix}
 0 \\
 \frac{1}{\sqrt2}\\
 \frac{1}{\sqrt2} 
\end{bmatrix}" , "\displaystyle e_{3}=\frac{X_{3}}{\mid\mid X_{3}\mid\mid}= \begin{bmatrix}
 0 \\
 \frac{1}{\sqrt2}\\
 -\frac{1}{\sqrt2} 
\end{bmatrix}"


Therefore,

The normalized model matrix is

"\displaystyle \bar{p}=\begin{bmatrix}
 1 & 0 & 0 \\
 0 & \frac{1}{\sqrt{2}} &\frac{1}{\sqrt{2}} \\ 
 0 & \frac{1}{\sqrt{2}} & -\frac{1}{\sqrt{2}}
\end{bmatrix}" and "\bar{p}^{-1}=\bar{p}^{T}=\frac{1}{\sqrt{2}}\displaystyle \begin{bmatrix}
 \sqrt{2} & 0 & 0 \\
 0 & 1 & 1 \\ 
 0 & 1 & - 1
\end{bmatrix}"


"\bar{p}^{-1}A\bar{p}= \frac{1}{\sqrt{2}}\displaystyle \begin{bmatrix}
 \sqrt{2} & 0 & 0 \\
 0 & 1 & 1 \\ 
 0 & 1 & - 1
\end{bmatrix}\begin{bmatrix}
 \ 1 & 0 & 0 \\
 0 & 3 & -1 \\ 
 0 & -1 & 3
\end{bmatrix}\begin{bmatrix}
 1 & 0 & 0 \\
 0 & \frac{1}{\sqrt{2}} &\frac{1}{\sqrt{2}} \\ 
 0 & \frac{1}{\sqrt{2}} & -\frac{1}{\sqrt{2}}
\end{bmatrix}"


"=\begin{bmatrix}
 \ 1 & 0 & 0 \\
 0 & 1 & 1 \\ 
 0 & 1 & -1
\end{bmatrix}=D"


We know that the quadratic form is "Q=X^{T}AX" where "X=\bar{P}Y"

then,

"Q=(\bar{P}Y)A(\bar{P}Y)"

"=(Y^{T}\bar{P}^{T})A(\bar{P}Y)\\=Y^{T}(P^{T}A\bar{P})Y\\=Y^{T}(\bar{P}^{-1}A\bar{P})Y\hspace{0.3cmsince\hspace{0.3cm}}\bar{P}^{T}=\bar{P}^{-1}\\=Y^{T}DY"

Therefore, the canonical form

"=\begin{bmatrix}
 y_{1} & y_{2} &y_{3} \\
 
\end{bmatrix} \begin{bmatrix}
 \ 1 & 0 & 0 \\
 0 & 2 & 0 \\ 
 0 & 0 & 4
\end{bmatrix} \begin{bmatrix}
 \ y_{1} \\
 y_{2} \\ 
 y_{3} 
\end{bmatrix}"


"= y_{1}^{2}+2y_{2}^{2}+4y_{3}^{2}"

which gives the canonical form.